# Portable, Massively Parallel Implementation of a Material Point Method for Compressible Flows

### v1

@@ 4.2 Sensitivity to numerical parameters
Having addressed the applicability of the method to high speed, compressible gas dynamics, we focus on the cylinder test case and test the dependence of the results on key numerical parameters. Specifically, we vary the number of grid cells, $N_{C}$, and the average number of particles per cell, $ppc$, and evaluate the effect on numerical stability and shock-capturing ability.
First, sufficiently high spatial resolution must be used to resolve shocks, otherwise the scheme becomes too diffusive. For the cylinder test case, testing evidence suggests that this minimum resolution amounts to around 3500 cells. Besides resolution, supersonic compressible fluid dynamics requires further care in setting the average number of particles per cell. Indeed, $ppc$ affects stability in a twofold fashion. First, for low $ppc$ values (e.g.,1), widespread oscillations are observed all along the simulation, due to almost randomly distributed empty cells. In a recent study, a $ppc$ value of 4 was found to be optimal for convergence of gas dynamics tests at lower speed [44], and, indeed, a $ppc=4$ solves this issue. However, for supersonic flow past a static obstacle, we observed that using $ppc=4$ might cause the formation of domain regions of empty or quasi-empty cells, which gets evident as grid resolution increases, during the initial transient phase from uniform, undisturbed conditions, as can be seen from Figure 6.
@@
Such emptying causes later formation of nonphysical cavitation-like effects during the transient that precedes equilibrium, which alter the dynamics, and, if particularly strong, can cause also spurious shock waves, that interact with the physical ones. These events can be prevented by increasing the average particles per cell number: at Mach 3, spurious shock waves become less relevant with $ppc=9$, and undetectable with $ppc=16$.
These effects, and in particular strong disruptions, are of no interest here, since the initial transient phase is non-physical, and derives only from the choice to start the simulations with undisturbed inflow conditions and particles position sampled from a uniform random distribution instead that from a previously computed equilibrium condition, e.g. the result of an already converged simulation. However, it could be a relevant aspect if fast changing dynamic situations have to be simulated, for example with a moving cylinder in a FSI framework, since they would alter its dynamics. In such cases, a sufficient number of particles per cell should be used. By contrast, a $ppc=4$ value can be employed when simulating equilibrium conditions: numerical experiments show that equilibrium is reached notwithstanding spurious instabilities in the initial transient phase (see Figure 7 for a snapshot).
However, for linear grid basis functions, numerical oscillations are present for any $ppc$ value. Therefore, additional artificial viscosity is essential to compute numerical quantities, even though going high order might be necessary to properly resolve fine structures. For example, in the cylinder test case the standard deviation on pressure, temporally averaged over enough different equilibrium instants, can decrease by about one half with the addition of a proper artificial viscosity (Figure 7 shows also effects of numerical viscosity on a single snapshot).
Figure 7: Mach 3 flow past a cylinder test case, equilibrium snapshot with $N_{C}=10^{5}$, $ppc=4$ and no artificial viscosity (above), and $N_{C}=10^{6}$, $ppc=16$ and artificial viscosity. Despite the very different setting, which determines a different initial behaviour, the two simulations foresees according pressure around the cylinder at equilibrium, and only slight differences in the far rear.
Finally, a sufficient randomness of particles position, especially in particles recycling, is fundamental to obtain stable and physically meaningful simulations, in agreement with results in the literature on Particle-In-Cell methods in plasma physics. On longer runs (i.e. above tens of physical seconds) a symplectic time integration scheme might be useful [31].
@@ 4.3 Computational performance
The first finding is that both kernels scale linearly with the total number of particles, in agreement with their implementation (Figures 8 and 9), while pure grid kernels, such as boundary conditions enforcement, scale with the number of nodes. For this reason, we applied a global particles reordering step in the algorithm and studied its effects on performance. Indeed, the ordering is done so that particles which are in the first grid cell after moving are assigned the lowest global indices, followed by those in the second grid cell, and so on moving along the grid. Thus, data which will be accessed at the next time iteration will be ordered, with contiguous particles having contiguous indexes, keeping data locality across the simulation, which is of utmost importance on our target architecture. The reordering turns out to have a significant impact on the G2P kernels, resulting in a further 4x speed-up factor, while preserving the linear scaling (Figure 8).
@@
Portability
Finally, we test the application of the developed implementation also on a different architecture and compare the performance, addressing the performance portability of the MPM code. Specifically, we execute the cylinder test case on $242x$ 24 CPU cores, running the nine numerical settings described above, without enforcing particles re-ordering, varying the number of cells and of particles per cell and studying how this affects the main computational kernels (Figure 11).
Figure 11: Profiling of main computational kernels: total G2P and P2G kernels execution time is compared for the cylinder test case, varying $N_{C}$ and $ppc$, without re-ordering, on $242x$ 24 CPU cores.
While portability is automatically granted, CPU performance turns out to be worse than single-GPU performance, even though 2x24 CPU cores are employed in parallel. This is consistent with expectations that MPM is better suited to GPU architectures.
@@
  • the execution time required by all the other kernels considered together is considerably smaller than the time required by P2G and G2P ones. This confirm the computational relevance of P2G and G2P also on different hardware.
The main difference with respect to the GPU test lies in the comparison between P2G and G2P performance, with the former significantly slower than the latter on the CPU. This is most probably due to our choice of writing an ad-hoc atomicAdd function, based on std::atomic_ref, with the same signature of the more performant CUDA’s atomicAdd, to enhance portability and guarantee parallelism. An algorithm to directly avoid the data-race present in P2G-like kernels, and, thus, to provide a natively portable, parallel and atomic operation free implementation, is proposed in the next section.
@@ 5 Discussion and conclusion
This paper has showcased the potential of a performance portable implementation of the Material Point Method for compressible flows. Insight has been given on algorithm adaptations which will provide further benefits, moreover, the simulation of standard test cases of supersonic flow has provided results comparable with the literature and with established software runs with the same polynomial order (first order - Lagrangian hat functions) and number of degrees of freedom. These results show that MPM is able to simulate Mach 3 gas dynamics, giving qualitatively good results with a prototype developed modifying existing implementations [58, 12], and serve as further step towards the design of a highly efficient MPM-based supersonic FSI solver.
As known from the literature [57, 42] and as also noted in a previous implementation [2], $\mathbb{Q}1$ MPM suffers from strong numerical oscillation issues, especially in the vicinity of shocks. The issue is tackled in this paper by adding numerical viscosity in the form of additional terms in the momentum equation [56]. A more natural and accurate choice could be obtained by using quadratic B-Spline basis functions [49, 17], a code feature which is currently under development.
Performance tests on NVIDIA A100 GPUs have confirmed that the main computational steps of the MPM algorithm are the Particles-to-Grid and Grid-to-Particles kernels, and that they scale well with the size of the problem. Moreover, adding a reordering step is beneficial to data locality and, consequently, performance, while data races in the P2G kernel can be tackled CUDA’s atomics at acceptable cost, even though alternative exist. Last, this approach has shown to be easily portable to different hardware architectures, as CPUs.
@@
  • For the algorithm implementation, the use of CUDA atomic operation of GPUs choice was chosen for simplicity, also because it shown an acceptable impact on computational time and scalability (section 4). To reach the full parallelisation capabilities proved by the G2P kernels also for the P2G ones, P2G algorithm adaptations will be necessary. An attractive and and more portable route would involve avoiding atomics, and writing a grid colouring algorithm that can be implemented in an architecture-independent way, relying only on standard libraries functions. Its main steps are:
@@
  • On the High Performance Computing side, avenues of development include the porting of the existing code to different backends, such as HIP and ROCm, the support of multi-GPU and multi-node architectures via MPI, and a grid-particles adaptive refinement through a quad-tree approach [47, 13]. Moreover, currently the code employs Structure-of-Arrays for the memory, since, differently from arrays of arbitrary structures, it provides coalescing, which improves memory efficiency [27]. However, recent developments on GPUs computing literature [54, 20] show that the best performance is achieved by adopting an Array of SoA (AoSoA). When addressing the MPI-based multi-GPU implementation, this possibility could be explored: indeed, the current data structure gives a conceptually straightforward extension towards multi-GPU implementations, where a further level of parallelisation through an MPI-based domain decomposition will make each GPU hold a SoA as the current one, globally giving rise to an AoSoA data structure.
@@ Acknowledgements
Research was sponsored by the Leonardo Labs, Leonardo S.p.A., through the PhD scholarship “Multi-GPUs based PIC / MPM methods for Compressible Flows” funded at Politecnico di Milano and by ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data, and Quantum Computing funded by European Union - NextGenerationEU. Views and conclusions contained in this document are those of the author and should not be interpreted as representing the official policies, either expressed or implied, of Leonardo Labs or ICSC.

### v2

@@ 4.2 Sensitivity to numerical parameters
Artificial viscosity
Artificial viscosity introduced in section 2 is composed of two main terms: a quadratic one and a linear one. The first, introduced by von Neumann-Richtmyer [59], is a viscous term that, being added to the pressure in the vicinity of shock waves, reduces numerical oscillations and enables the capturing of the shock. The second, introduced by Landshoff [36], is a linear term added to the quadratic one to act specifically in the region behind the shock front. Further details on the application of this viscosity to MPM can be found in [63].
In this paragraph we study the effect of the artificial viscosity, comparing the final density obtained enforcing $c_{0}=1,c_{1}=1$ in equation (5) and shown in Figure 5, with the one obtained by:
  • turning off the artificial viscosity;
  • applying only the von Neumann-Richtmyer term, enhanced by a factor 10;
  • applying only the Landshoff term, enhanced by a factor 10.
Results are in agreement with theoretical expectations (Figure 11) and confirm the importance of this approach in the linear MPM scheme adopted here.
Figure 11: Sod Shock Tube - density trend at time $t=0.143$, varying artificial viscosity coefficients.
Resolution
Having addressed the applicability of the method to high-speed transonic and supersonic compressible gas dynamics, we focus on the cylinder test case and test the dependence of the results on key numerical parameters. Specifically, we vary the number of grid cells, $N_{C}$, and the average number of particles per cell, $ppc$, and evaluate the effect on the pressure of the flow.
First, sufficiently high spatial resolution must be used to actually show shock waves; indeed, notwithstanding the use of artificial viscosity as shock capturing technique, with $N_{C}\sim 10^{3}$ cells we are not able to see any shock. However, a run using $N_{C}\simeq 10^{4}$ cells is able to show flow features which are in agreement with what expected from the literature, as can be seen in Figure 12.
Besides resolution, supersonic compressible fluid dynamics requires further care in setting the average number of particles per cell. Indeed, $ppc$ affects stability in a twofold fashion.
First, for low $ppc$ values (e.g.,1), widespread oscillations are observed all along the simulation, due to almost randomly distributed empty cells. In a recent study, a $ppc$ value of 4 was found to be optimal for convergence of gas dynamics tests at lower speed [50], and, indeed, a $ppc=4$ solves this issue. However, for supersonic flow past a static obstacle, we observed that using $ppc=4$ might cause the formation of domain regions of empty or quasi-empty cells, which gets evident as grid resolution increases, during the initial transient phase from uniform, undisturbed conditions, as can be seen from Figure 13. Such emptying causes later formation of nonphysical cavitation-like effects during the transient that precedes equilibrium, which alter the dynamics, and, if particularly strong, can cause also spurious shock waves, that interact with the physical ones. These events can be prevented by increasing the average particles per cell number: at Mach 3, spurious shock waves become less relevant with $ppc=9$, and undetectable with $ppc=16$.
These effects, and in particular strong disruptions, are of no interest here, since the initial transient phase is non-physical, and derives only from the choice to start the simulations with undisturbed inflow conditions and particles position sampled from a uniform random distribution instead that from a previously computed equilibrium condition, e.g. the result of an already converged simulation. However, it could be a relevant aspect if fast changing dynamic situations have to be simulated, for example with a moving cylinder in a FSI framework, since they would alter its dynamics. In such cases, a sufficient number of particles per cell should be used.
By contrast, when interested in main equilibrium conditions, a $ppc=4$ value appears to provide sufficient solution quality. Numerical experiments show that equilibrium is reached notwithstanding spurious instabilities in the initial transient phase, as shown in temporal averaged pressure in Figure 14. However, as shown in the same figure, a higher resolution allows the accurate representation of small areas, such as the ones in the immediate rear of the cylinder, where quantities, such as the average Mach number, change in a non-negligible way.
Figure 12: Mach 3 flow past a cylinder test case with $N_{C}=10^{4}$, $ppc=4$, pressure field at $t\simeq 1s$.
@@
Figure 14: Mach 3 flow past a cylinder test case, at low resolution (top row, $N_{C}=10^{4}$, $ppc=4$) and high resolution (bottom row, $N_{C}=10^{6}$, $ppc=16$), from left to right: average pressure, pressure standard deviation (represented only to provide a clearer image of the shocks), average Mach number. Values computed as arithmetic average over 85 snapshots corresponding to physical simulation time from 15 seconds to 100 seconds.
@@ 4.3 Computational performance
The first finding is that both kernels scale linearly with the total number of particles, in agreement with their implementation (Figures 15 and 16), while pure grid kernels, such as boundary conditions enforcement, scale with the number of nodes. For this reason, we applied a global particles reordering step in the algorithm and studied its effects on performance. Indeed, the ordering is done so that particles which are in the first grid cell after moving are assigned the lowest global indices, followed by those in the second grid cell, and so on moving along the grid. Thus, data which will be accessed at the next time iteration will be ordered, with contiguous particles having contiguous indexes, keeping data locality across the simulation, which is of utmost importance on our target architecture. The reordering turns out to have a significant impact on the G2P kernels, resulting in a further 4x speed-up factor, while preserving the linear scaling (Figure 15).
noindent As for overall computational time, we note that its value largely depends on the requested resolution. On a NVIDIA A100 GPU, to simulate one second of gas flow around the cylinder with $4\cdot 10^{4}$ particles requires about one second of computational time on the device, with $16\cdot 10^{6}$ particles it requires about $3$ minutes. This is reasonably fast, especially considering that the latter resolution is rather excessive for the 2D cylinder test case. For comparison, 3D snow avalanches were recently simulated via MPM with $10$ to $30$ million particles, see [25].
From a previous study [2], we expect that even better performance is achievable. However, with this work we are mostly interested in investigating the potential of MPM when simulating compressible fluids on various GPU accelerated architecture, we thus prioritise portability over efficiency, to be able to compare exactly the same code on different architecture. Portability has been reached also by means of a higher-level approach with respect to low-level CUDA-C, which introduces some level of indirection, with libraries as Thrust that rely on managed memory.
In the next paragraph, we exploit the flexibility granted from this approach to test the code on different hardware architectures, both on CPUs and GPUs.
@@
4.4 Portability
We test the application of the developed implementation on different hardware architectures and compare the performance, addressing the portability of the MPM code. Specifically, we execute the cylinder test case at several resolutions, on different Intel and AMD CPUs and NVIDIA and AMD GPUs, with the same setting described in the previous sections, without enforcing particles re-ordering.
For the first test, we vary the number of cells and of particles per cell and study how this affects the main computational kernels on an AMD CPU, employing $242x$ 24 cores in parallel (Figure 18).
While portability is automatically granted, CPU performance turns out to be worse than single-GPU performance, even though $242x$ 24 CPU cores are employed in parallel. This is consistent with expectations that MPM is better suited to GPU architectures.
@@ 4.4 Portability
  • the execution time required by all the other kernels considered together is considerably smaller than the time required by P2G and G2P ones. This confirms the computational relevance of P2G and G2P also on different hardware.
The main difference with respect to the GPU test lies in the comparison between P2G and G2P performance, with the former significantly slower than the latter on the CPU. This is most probably due to our choice of writing an ad-hoc atomicAdd function, based on std::atomic_ref, with the same signature of the more performing CUDA’s atomicAdd, to enhance portability and guarantee parallelism. An algorithm to directly avoid the data-race present in P2G-like kernels, and, thus, to provide a natively portable, parallel and atomic operation free implementation, is proposed in the next section.
Figure 18: Profiling of main computational kernels: total G2P and P2G kernels execution time is compared for the cylinder test case, varying $N_{C}$ and $ppc$, without re-ordering, on $242x$ 24 CPU cores (AMD 7402 CPU).
For the second test, we focus on the heavier test case, the cylinder with $16$ million particles, and simulate one physical second on different architectures, comparing the total wall clock time, referred to the time required by the device, being either a CPU or a GPU; see Figure 19. It is interesting to note that the chosen implementation approach allows the porting of the code to AMD GPUs without any modification.
Figure 19: Comparison of performance on different architectures: overall computational time to simulate 1 physical second of the cylinder test case with $N_{C}=10^{6}$, $ppc=16$.
Lastly, we focus on the comparison between the two GPU architectures.
We can notice that at low computational load, performance is basically equivalent, especially if the higher degree of code optimisation allowed by the NVIDIA compilers is taken into account, even though the AMD GPU seems to be more sensible to the kernels with atomicAdd; see Figure 20.
However, at heavier loads, despite the comparable specs, the AMD MI210 starts to display longer wallclock times than the NVIDIA A100, see Figure 21 for a comparison of the overall computational time on the device. This difference might be due to different aspects, such as having developed the code, and taken performance based design choices, only on NVIDIA A100 GPUs, and having ported it to different hardware afterwards. However, GPU performance is overall competitive with respect to CPU. For example, the heavier test case is 7 to 8 times slower on MI210 than on A100, but anyway 3 to 4 times faster than parallel execution via Intel OneAPI-TBB on Intel Xeon with 2x24 cores.
Most importantly, these results enable us to confirm the portability of the approach.
Figure 20: Comparison of performance on different GPUs: analysis of the average computational time for a low resolution test case $N_{C}=10^{4}$, $ppc=4$, corresponding to about half computational second to simulate 1 physical second.
Figure 21: Comparison of performance on different GPUs: overall computational time required for $Nt=10^{5}$ time loop iterations, varying resolution from $N_{C}=10^{4}$, $ppc=4$ to $N_{C}=10^{6}$, $ppc=16$.
@@ 5 Discussion and conclusion
This paper has showcased the potential of a massively parallel, portable implementation of the Material Point Method for compressible flows. Insight has been given on algorithm adaptations which will provide further benefits, moreover, the simulation of standard test cases of compressible flow has provided results comparable with the literature and with established software runs with the same polynomial order (first order shape functions - Lagrangian hat functions) and number of degrees of freedom. These results show that MPM is able to simulate compressible gas dynamics, giving qualitatively good results with a prototype developed modifying existing implementations [65, 14], and serve as further step towards the design of a highly efficient MPM-based transonic and supersonic FSI solver.
As known from the literature [64, 48] and as also noted in a previous implementation [2], $\mathbb{Q}1$ MPM suffers from strong numerical oscillation issues, especially in the vicinity of shocks. The issue is tackled in this paper by adding numerical viscosity in the form of additional terms in the pressure field [63]. A more natural and accurate choice could be obtained by using quadratic B-Spline basis functions [55, 19], a code feature which is currently under development [^8] https://github.com/carlodefalco/quadgrid.
Performance tests on NVIDIA A100 GPUs have confirmed that the main computational steps of the MPM algorithm are the Particles-to-Grid and Grid-to-Particles kernels, and that they scale well with the size of the problem. Moreover, adding a reordering step is beneficial to data locality and, consequently, performance, while data races in the P2G kernel can be tackled CUDA’s atomics at acceptable cost, even though alternative exist. Last, this approach has shown to be easily portable to different hardware architectures, as different vendors GPUs and CPUs.
@@
    A 3D extension will facilitate a further growth in computational load, but this will not introduce conceptual main changes in the code, also thanks to the high level, portable approach, which does not refer explicitly to specific, small size memories. However, it could justify a more efficient I/O, e.g., moving from JSON to HDF5. Simulations characterised by a higher computational load would also further benefit from a future multi-GPUs approach.
    The upgrade to a higher order MPM via proper shape function, as B-Splines, would not require a significant change in the numerical method, but will introduce some challenges linked to parallelism.
    In particular, stencil growth due to expansion of shape function support to neighbouring grid elements will introduce halo overlapping, leading to an increase in communication costs, as well as to a constraint on the maximum level of achievable parallel granularity.
    Few works on this topic exist, such as [61], and, since they simulate continuum mechanics materials instead of compressible gases, they can benefit from sparse data structures. However, it is already possible to anticipate that shape function regularity in the current MPM formulation could grow efficiently only up to a certain maximum, such as quadratic or cubic B-Spline, and start to have an impact on performance as shape function support expands.
  • For the algorithm implementation, the use of atomic operations to handle data races was chosen for simplicity, also because it showed an acceptable impact on computational time and scalability on NVIDIA A100 GPUs and a simple portability to other architectures (section 4). To reach the full parallelisation capabilities proved by the G2P kernels also for the P2G ones, P2G algorithm adaptations will be necessary. An attractive and and more portable route would involve avoiding atomics, and writing a grid colouring algorithm that can be implemented in an architecture-independent way, relying only on standard libraries functions. Its main steps are:
@@
  • On the High Performance Computing side, avenues of development include the support of multi-GPU and multi-node architectures via MPI, and a grid-particles adaptive refinement through a quad-tree approach [53, 15]. Moreover, currently the code employs Structure-of-Arrays for the memory, since, differently from arrays of arbitrary structures, it provides coalescing, which improves memory efficiency [30]. However, recent developments on GPUs computing literature [61, 22] show that the best performance is achieved by adopting an Array of SoA (AoSoA). When addressing the MPI-based multi-GPU implementation, this possibility could be explored: indeed, the current data structure gives a conceptually straightforward extension towards multi-GPU implementations, where a further level of parallelisation through an MPI-based domain decomposition will make each GPU hold a SoA as the current one, globally giving rise to an AoSoA data structure.
@@ Acknowledgements
Research was sponsored by the Leonardo Labs, Leonardo S.p.A., through the PhD scholarship “Multi-GPUs based PIC / MPM methods for Compressible Flows” funded at Politecnico di Milano and by ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data, and Quantum Computing funded by European Union - NextGenerationEU. The simulations discussed in this work were, in part, performed on the HPC Cluster of the Department of Mathematics of Politecnico di Milano which was funded by MUR grant Dipartimento di Eccellenza 2023-2027. C. de Falco and P. J. Baioni are members of the GNCS group of INdAM. Views and conclusions contained in this document are those of the author and should not be interpreted as representing the official policies, either expressed or implied, of Leonardo Labs, ICSC, MUR or GNCS-INdAM.